OpenStudy (devincoolage1001):

−3(x + 3) = −3(x + 1) − 5 solve for x

OpenStudy (devincoolage1001):

6 −6 All real numbers No solution

OpenStudy (amoghk):

−3(x + 3) = −3(x + 1) − 5 -3x -9 = -3x - 3 - 5 -3x - 9 = -3x - 8 9=8 (which is wrong) Hence no solution

OpenStudy (devincoolage1001):

Thank you.